OpenStudy (anonymous):

Solve 5x – 6y = –38 3x + 4y = 0 (4, 3) (–4, 3) (4, –3) (–4, –3)

OpenStudy (anonymous):

x=-4y/3 thus -20y/3-6y=-38 -38y/3=-38 y=3 x=-4 (-4,3)